The sum of the measures of the interior angles of a convex polygon is 1980. How many sides does the polygon have?
Answer by stanbon(75887)

The sum of the measures of the interior angles of a convex polygon is 1980. How many sides does the polygon have?
---------------------------
Fact: The sum of the angles of any regular n-gon is (n-2)180 degrees.
-------------
Your Problem:
(n-2)180 = 1980
(n-2) = 11
n = 13 sides
--------------------
Cheers,
Stan H.